What is the molecular formula of the compound if its empirical formula is and its molecular weight is 90.0 amu


1
Expert's answer
2021-10-11T06:10:09-0400

Emperical formula weight of CH

2

O is 30 (12+2+16)

Molecular formula weight is 90

So,

n= 90/30= 3

Molecular formula=3×CH2O

Hence molecular formula is C3H6O3